The order is for heparin 300 units / hr. The pharmacy sends an IV of 1000 mL D5W with 10,000 units of heparin. How many mL / hr should the nurse set on the IV pump? ___________ mL / hr
Answer #1

Given concentration of heparin = 10,000 unit per 1,000 ml of D5W

  • Therefore 1 unit of heparin will be present in = 1000/10,000 ml of D5W = 0.1 ml
  • Or 0.1 ml of D5W contains 1 unit of heparin
  • Therefore 300 unit of heparin will be present in 300 × 0.1 = 30 ml of D5W

Rate of IV = 30 ml/hr

Feel free to comment if you have any problem in understanding this answer.